ASPEARS:

Envision Algebra 2 lesson 4.5 and 5.1 practice problem solving

Rylee88:

Make sure to provide more information, graphs, or pictures of this so people can help please!

ASPEARS:

okay

ASPEARS:

the practice number 15-30

ASPEARS:

numbers 25-43

Joe348:

So for number 15 you have to plug in a number for x to figure out if it equals ten \[\frac{ 1 }{ x-3 }=10\] I would start from 1-10 to see if one of those fit. Once you plug these numbers in for x, you realize that none of them equals 10, you find the closest number to equal ten. \[\frac{ 1 }{ 3.1-3}=10\] \[x=3.1\]

Joe348:

For 16-18, you would do the same thing as #15. Plug in a number for x.